# Few-Shot API Attack Anomaly Detection in a Classification-by-Retrieval Framework

### v1

@@ IV-A Data Pre-Processing
For both training and inferring the detection model, we employ a unique pre-processing technique consisting of three phases, as depicted in Figure 1 in steps 2.1 and 3.1. First, decoding URL special characters, decompressing request body content, and converting every character within the request data string into lowercase. Then, validating request headers to ensure they are formatted correctly and extracting the endpoint definition, as depicted in Figure 1 2.1.1 and 3.1.1. API request headers typically provide information about the request context, supply authentication credentials, and provide information about the client (e.g., a person, a computing device, and/or a browser application) that had initiated the API  [39]. API request header fields are typically derived from a limited set of options. Accordingly, the data preparation also uses a fixed set of rules to validate the content of request headers and filter-out headers according to these rules. Headers that include valid or approved strings may be transferred to their destination as an API request and may be excluded from additional processing. For example, request headers may include host strings, which specify host or Internet Protocol (IP) addresses and/or port numbers of a server to which the API request is being sent. Valid IPv4 syntax should be in the format of: $(0\leq n<256).(0\leq n<256).(0\leq n<256).(0\leq n<256):(1\leq n\leq 65535)$.
@@
Lastly, as shown in Figure 1 2.1.2 and 3.1.2, we convert received requests into abstracted versions based on a conversion schema. For instance, we replace non-numeric single characters with the string ”chr”, which serves as a representative, abstract version of the original request string. Another example involves converting non-textual symbols into predefined textual strings. for instance, colons (”:”) may be converted to the string ”colon”. During the pre-processing stage, the API language has been refined to achieve optimally, enabling it to describe various transactions in a consistent format without sacrificing their original meanings. In fact, in the majority of cases, these requests have been intelligently merged into a single text representation. This consolidation not only streamlines the data but also ensures that the essential information pertaining to different transactions remains preserved.
@@
IV-B Unsupervised ML Language Model
Our framework leverages FastText for unsupervised learning. FastText possesses the capability to encapsulate substantive knowledge about words while integrating morphology details, a crucial aspect for API attack detection. While deep learning models have excelled state-of-the-art results across various NLP tasks, to the best of our knowledge, no previous NLP pre-trained model on the API traffic domain has been publicly published. In the training phase of the method, we built a single generic FastText language model (based on  [29]) from scratch using the normal API traffic collected by crawling Tranco’s list of the most popular websites, as can be seen in Figure 1 Phase 1. For training the generic language model, we used the default hyper-parameters of  [29], which encompass a learning rate of 5%, a word vector size of 100, and a context window size of 5. The model was then trained for 5 epochs. We utilized the CBOW model as the dataset is relatively large, and CBOW embeddings are precise enough for anomaly detection and computed in a shorter time than skip-gram [40]. For training the detection model and inferencing, we extract the vector representation of words for every input line.
IV-C ANN
We obtained ANN to identify normal representation of an API endpoint. We train a single detection model, which is used to describe a normal representation of all endpoints. During the detection model training stage, each API request is represented as a vector in the textual embedding space, including endpoint information. We employ cosine distance to measure the similarity between data points as it has been applied in numerous text mining endeavors, such as text classification, and information retrieval  [41]. Additionally, it has been proven to be effective for Out-of-Distribution (OOD) detection tasks  [42, 43]. Cosine similarity is a widely used measure of similarity that calculates the angle formed by a pair of vectors. When measuring the similarity between two patterns, the Euclidean distance increases as they become less similar, while the cosine similarity increases as they become more similar. Unlike Euclidean distance, cosine similarity is unaffected by the magnitude of the vectors being compared [44]. The embedding vector feeds the Hierarchical Navigable Small Worlds (HNSW) graph [45] to build new indexes of data points. During the detection stage, the model compares new API vectors of incoming API requests in relation to API vectors of the same API endpoint information to evaluate normality or anomaly of the incoming API request. The comparison is made by querying the similarity between the input vector to the closest k objects.
@@ IV-C ANN
V Experimental Design
@@
API security papers are not as prevalent as the technology itself, despite being one of the most influential technologies  [46]. This is particularly evident in the scarcity of ready-to-use publicly available API datasets  [47]. Several of these datasets are obsolete and unsuitable for usage, with some lacking traffic diversity and volumes, and others failing to encompass a variety of attacks, such as ECML/PKDD 2007  [48] and CSE-CIC-IDS 2018 [49].Consequently, researchers resort to creating customized datasets by primarily utilizing open-source vulnerable web applications like DVWA, BWAPP, and Mutillidae, and employ automated penetration tools such as SQLMAP [50], SQLNINJA [51], and OWASP ZAP [52] to gather malicious payloads  [53]. Therefore, this research is evaluated on two datasets: CSIC 2010  [54] and ATRDF 2023  [55]. The HTTP CSIC 2010 dataset  [54] is widely used  [56, 33, 57, 23, 58, 59, 27, 24, 25, 26, 60, 61, 62] in the field of malicious web traffic detection. This dataset was created by the Spanish Research National Council (CSIC). It is a sample of the traffic occurring on the Spanish e-commerce web application. The dataset includes attacks such as SQL injection (Figure 4(a)), buffer overflow, information gathering, files disclosure, CRLF Injection, XSS, static attacks and unintentional illegal requests. While unintentional illegal requests lack malicious intent, they deviate from the typical behavior of the web application and exhibit a different structure compared to regular parameter values. For instance, as shown in Figure 4(b), an invalid DNI (Spanish national ID number) was marked as an anomaly. We divided the dataset into two segments: the training portion, which included 36,000 normal requests and was exclusively utilized for representation learning, and the inference portion, which comprised both 36,000 normal and 25,000 anomalous traffic that was encoded by the model and employed for detection. It has 38 different endpoints, 8 of which have no normal representation and were excluded from our experiment.
@@
The API Traffic Research Dataset Framework (ATRDF) [55] is a recently published HTTP dataset publicly available which includes 18 different API endpoints. The dataset includes attacks such as Directory Traversal, Cookie Injection, Log4j (Figure 4(c)), RCE, Log Forging, SQLi, and XSS. The dataset contains 54,$0$ normal and 78,000 abnormal sets of request and response.
@@
In response to the unavailability of a publicly accessible pre-trained model specialized for the API domain, a generic language model was developed to establish a reliable baseline for various detection models. However, developing a robust language model necessitates a significant amount of training data. To address this requirement, an extensive dataset of 1,061,095 API examples was collected. This dataset was obtained by performing a comprehensive data collection process, involving the crawling of a random sample of websites from the Tranco top websites list [^2] https://tranco-list.eu/. The Tranco list is an invaluable resource for cyber-security research as it provides a publicly available compilation of the top one million most popular domains, ranked based on a combination of four reputable lists: Alexa, Cisco Umbrella, Majestic, and Quantcast. Additionally, the Tranco list offers the advantage of being able to filter out unavailable or malicious domains, making it a valuable asset for our research [63].
@@
Figure 4: An example of anomaly requests from CSIC 2010 and ATRDF 2023 datasets where endpoint definition is marked in green, abnormal payload is marked in red.
VI Evaluation Metrics
In the realm of anomaly detection, it is commonplace to integrate a binary classification layer into the model architecture. This is due to the fundamental objective of distinguishing between normal and abnormal instances. To evaluate our architecture, we suggest using several performance measures in various experiments, including precision, recall, accuracy and F1-score. Actual values are represented as True and False by (1) and (0), respectively and predicted as Positive and Negative values by (1) and (0), respectively. Predicted possibilities of classification models are obtained through the expressions TP, TN, FP, and FN.
@@ VI Evaluation Metrics
In the current problem, two classes are represented by Positive and Negative, where the positive class corresponds to an abnormal API request and the negative class corresponds to a normal API request.
VII Experimental Results
To understand better the two datasets, we use the t-Distributed Stochastic Neighbor Embedding (T-SNE) method for dimensionality reduction to graphically depict our high-dimensional datasets. The plots in Figure 5 indicate distinct separation among classes within a reduced dimensional space. This suggests that requests with similarities tend to be grouped together, enabling the exploration of a neighborhood for any given sample. Close records from different classes suggest that certain requests are quite similar, causing the embedding to overlap between the two classes. The result for ATRDF 2023 shows no overlap and clear separation between the classes while for CSIC 2010, some records from the two classes were found to be similar. We identified that most of the anomalous requests which overlap with normal requests actually have no malicious payload and are categorized as unintentional illegal requests.
@@ VII Experimental Results
Figure 5: Vector representations reduced to 2D with t-SNE
@@
Then, we compare our method with fourteen detecting outlying objects in multivariate data baseline models [64]: Feature Bagging (FB), Histogram-based Outlier Detection (HBOS), Isolation Forest (IF), Local Outlier Factor (LOF), Minimum Covariance Determinant (MCD), One-class SVM (OCSVM), Principal Component Analysis (PCA), Copula-Based Outlier Detector (COPOD), Deep One-Class Classification for outlier detection (DeepSVDD), Clustering-Based Local Outlier Factor (CBLOF), Outlier detection based on Gaussian Mixture Model (GMM), Kernel Density Estimation (KDE), Linear Model Deviation-based Outlier Detection (LMDD), Quasi-Monte Carlo Discrepancy Outlier Detection (QMCD). We apply default hyper-parameters as provided by the original source code across all models for consistency and use the same pre-processed dataset to facilitate optimal comparison.
@@
Furthermore, we conducted an ANN benchmark to effectively contrast our framework with other nearest neighbor algorithms. This benchmark involved generating algorithm instances based on configuration file written in YAML format that defines the different methods and algorithms. At the top level, the point type is specified, followed by the distance metric, and finally, each algorithm implementation to be evaluated. Each implementation specifies the Python library, and additional entries provide the necessary arguments. For clarity, an illustrative example of this configuration file is presented in Figure 6, while the complete file is available in the project GitHub repository  [^3] https://github.com/ArielCyber/FT-ANN-Journal/blob/main/ann_benchmark.yaml. Both the ”space” and ”run_groups” lists encompass arguments that should be included at the beginning of every invocation. Each algorithm defines one or more ”run groups,” each of which is expanded into several lists of arguments. The Cartesian product of these entries results in numerous argument lists. For instance, consider the hnswlib entry depicted in Figure  6. This expands into three distinct algorithm instances: Cosine (Cosine Similarity), L2 (Squared L2), and IP (Inner product). Each of these instances undergoes training before being utilized for various experiments. Initially, experiments are conducted with different values of k, representing the number of neighbors to return (e.g., [10, 50, 100, 300, 400, 500, 1000, 2000, 2500, 3000]). Subsequently, experiments are conducted with varying ef_construction values, which denote the size of the dynamic list used during index construction. A larger ef_construction value indicates a higher quality index but also results in longer build times (e.g., [10, 20, 40, 80, 120, 200, 400, 600, 800]). Throughout each run, pertinent information is recorded, including the algorithm name, the time taken to construct the data structure used for indexing, and the outcomes of every query. These query outcomes encompass the neighboring points returned by the algorithm, the duration required to locate these neighbors, and the proximity between the neighbors and the query point. We leveraged the query results to compute the corresponding confusion matrix, enabling us to thoroughly evaluate the classification-by-retrieval performance of each algorithm. We conducted our benchmarking analysis by assessing the performance of various algorithm implementations, including Nmslib, Hnswlib, Bruteforce Blas, Balltree, KDtree, CKDtree, Annoy, Faiss, and RPForest, all of which were evaluated using the publicly accessible ANN-Benchmarks tool as a reference framework [65]. In order to gain a deeper comprehension of how the embedding layer influences the detection outcome, we assessed all anomaly detection baseline models and the ANN benchmark using two additional prominent language models: BERT [66], and RoBERTa [67]. Each model was individually trained from scratch and subsequently subjected to evaluation. We employed the RoBERTa model, specifically the RoBERTaForMaskedLM class, with a language modeling head on top [^4] RoBERTa implementation by Hugging Face - https://huggingface.co/docs/transformers/model_doc/RoBERTa. This model was trained with a maximum sequence length of 512, utilizing 12 hidden layers and 12 attention heads. The training process spanned 10 epochs, employing a batch size of 16, which aligns with a similar approach in a prior study by  [23]. Likewise, we employed the BERT model, specifically BertForMaskedLM, which also incorporates a language modeling head on top [^5] BERT implementation by Hugging Face - https://huggingface.co/docs/transformers/model_doc/bert. This model underwent training with a maximum sequence length of 512, utilizing 4 hidden layers and 4 attention heads, as was the case in a similar task outlined in  [68].
@@
We used Intel(R) Xeon(R) Silver 4214R CPU @ 2.40GHz to evaluate the effectiveness of each technique. Each endpoint was evaluated separately, and the average score for all endpoints was used to measure overall detector performance. We evaluate model performance by measuring both execution time and F1-score during training and testing. Training time is usually longer than testing time as it entails parameter optimization, a computationally intensive task. Conversely, testing time is relatively faster as it only involves applying the pre-trained model to new data and predicting their likelihood of being outliers. It is worth noting that detecting anomalies in real-time is critical in mitigating the effects of an attack. Therefore, it is advisable to analyze each phase separately to better understand each model’s performance. Additionally, we aim to achieve better results compared to previous semi-supervised and unsupervised studies. Supervised studies should not be compared since they rely on labeled data.
@@
The results for the CSIC 2010 dataset, as can be seen from Table II, comparing against the conventional outlier detection baselines, NMSLIB-COS (FT-ANN) and PCA demonstrate the shortest times, with 0.0552 and 0.0575 seconds respectively, whereas AUTOENCODER and LMDD show significantly longer train times of 75.0644 and 225.8 seconds respectively. NMSLIB-COS achieves the highest F1 score of 0.9713, indicating its balanced precision-recall performance. Similarly, in terms of precision and recall, NMSLIB-COS again outperforms other algorithms with scores of 0.9538 and 0.9954 respectively. LMDD appears to exhibit the comparatively weakest performance across multiple metrics. With an F1 score of 0.3017, a precision of 0.528, and a recall of 0.2293, LMDD lags behind the other algorithms. Furthermore, its training time of 225.8 seconds is considerably longer than most other methods. The poor performance and unfavorable tradeoff between training time and results for LMDD could be attributed to its reliance on a dissimilarity function that may not be well-suited to the complex and diverse anomalies present [69].

### v2

@@ 3.1 Data Pre-Processing
For both training and inferring the detection model, we employ a unique pre-processing technique consisting of three phases, as depicted in Figure 1 in steps 2.1 and 3.1. First, decoding URL special characters, decompressing request body content, and converting every character within the request data string into lowercase. Then, validating request headers to ensure they are formatted correctly and extracting the endpoint definition, as depicted in Figure 1 2.1.1 and 3.1.1. API request headers typically provide information about the request context, supply authentication credentials, and provide information about the client (e.g., a person, a computing device, and/or a browser application) that had initiated the API  (Buchanan et al., 2018). API request header fields are typically derived from a limited set of options. Accordingly, the data preparation also uses a fixed set of rules to validate the content of request headers and filter-out headers according to these rules. Headers that include valid or approved strings may be transferred to their destination as an API request and may be excluded from additional processing. For example, request headers may include host strings, which specify host or Internet Protocol (IP) addresses and/or port numbers of a server to which the API request is being sent. Valid IPv4 syntax should have the following format: $(0\leq n<256).(0\leq n<256).(0\leq n<256).(0\leq n<256):(1\leq n\leq 65535)$.
@@
Lastly, as shown in Figure 1 2.1.2 and 3.1.2, we convert received requests into abstracted versions based on a conversion schema. For instance, we replace non-numeric single characters with the string "chr", which serves as a representative, abstract version of the original request string. Another example involves converting non-textual symbols into pre-defined textual strings. For instance, colons (":") may be converted to the string "colon". During the pre-processing stage, the API language has been refined to achieve optimally, enabling it to describe various transactions in a consistent format without sacrificing their original meanings. In fact, in most cases, these requests have been intelligently merged into a single text representation. This consolidation not only streamlines the data but also ensures that the essential information pertaining to different transactions remains preserved.
@@
3.2 Unsupervised ML Language Model
Our framework leverages FastText for unsupervised learning. FastText possesses the capability to encapsulate substantive knowledge about words while integrating morphology details, a crucial aspect of API attack detection. While deep learning models have excelled and commonly provide state-of-the-art results across various NLP tasks, to the best of our knowledge, no previous NLP pre-trained model on the API traffic domain has been publicly published. In the training phase of the method, we built a single generic FastText language model (based on  (Research, 2017)) from scratch using the normal API traffic collected by crawling Tranco’s list of the most popular websites, as can be seen in Figure 1 Phase 1. For training the generic language model, we used the default hyper-parameters of  (Research, 2017), which encompass a learning rate of 5%, a word vector size of 100, and a context window size of 5. The model was then trained for 5 epochs. We utilized the CBOW model as the dataset is relatively large, and CBOW embeddings are precise enough for anomaly detection and computed in a shorter time than skip-gram (Joulin et al., 2016). For training the detection model and inferencing, we extract the vector representation of words for every input line.
3.3 ANN
We obtained ANN to identify the normal representation of an API endpoint. We train a single detection model, which is used to describe a normal representation of all endpoints. During the detection model training stage, each API request is represented as a vector in the textual embedding space, including endpoint information. We employ cosine distance to measure the similarity between data points as it has been applied in numerous text mining endeavors, such as text classification, and information retrieval  (Li and Han, 2013). Additionally, it has been proven to be effective for Out-of-Distribution (OOD) detection tasks  (Techapanurak et al., 2020; Hsu et al., 2020). Cosine similarity is a widely used measure of similarity that calculates the angle formed by a pair of vectors. When measuring the similarity between two patterns, the Euclidean distance increases as they become less similar, while the cosine similarity increases as they become more similar. Unlike Euclidean distance, cosine similarity is unaffected by the magnitude of the vectors being compared (Xia et al., 2015). The embedding vector feeds the Hierarchical Navigable Small Worlds (HNSW) graph (Naidan et al., 2014) to build new indexes of data points. During the detection stage, the model compares new API vectors of incoming API requests with API vectors of the same API endpoint information to evaluate the normality or anomaly of the incoming API request. The comparison is made by querying the similarity between the input vector to the closest k objects.
@@ 3.3 ANN
4 Experimental Design
API security papers are not as prevalent as the technology itself, despite being one of the most influential technologies (Stone et al., 2022). This is particularly evident in the scarcity of ready-to-use publicly available API datasets (Torrano-Gimenez et al., 2011). For this research, we used HTTP datasets which reflect REST API traffic. REST API interactions are typically carried out over HTTP, and the resulting HTTP requests and responses are logged by the web server (Pautasso and Wilde, 2010). Several of these datasets, such as ECML/PKDD 2007 (Ware, 2007), are outdated and unsuitable for usage, with some lacking traffic diversity and volume, and others failing to encompass a variety of attacks. Similarly, the CSE-CIC-IDS 2018 (UNB, 2018) dataset contains only a small portion of web attacks and directs normal traffic requests to various web applications. Additionally, it derives anomalous traffic from DVWA application routes, complicating endpoint isolation, and often lacks textual payloads, making it unsuitable for our research. Consequently, researchers resort to creating customized datasets by primarily utilizing open-source vulnerable web applications like DVWA, BWAPP, and Mutillidae, and employ automated penetration tools such as SQLMAP (Damele and Stampar, ), SQLNINJA (Icesurfer and Nico, ), and OWASP ZAP (Bennetts, 2013) to gather malicious payloads, including the Fwaf project (Faizan, 2017).
@@ 4 Experimental Design
Therefore, this research is evaluated on two datasets: CSIC 2010 (Giménez et al., 2010) and ATRDF 2023  (Lavian et al., 2023). The HTTP CSIC 2010 dataset  (Giménez et al., 2010) is widely used  (Sánchez et al., 2021; Seyyar et al., 2022; Dawadi et al., 2023; Gniewkowski et al., 2021; Mac et al., 2018; Wang et al., 2018; Moradi Vartouni et al., 2019; Jemal et al., 2020; Niu and Li, 2020; Yu et al., 2020; Ito and Iyatomi, 2018; Jemal et al., 2021b; Yan and Xiong, 2020) in the field of malicious web traffic detection. This dataset was created by the Spanish Research National Council (CSIC). It is a sample of the traffic occurring on the Spanish e-commerce web application. The dataset includes attacks such as SQLi (Figure 4(a)), buffer overflow, information gathering, file disclosure, CRLF Injection, XSS, static attacks, and unintentional illegal requests. While unintentional illegal requests lack malicious intent, they deviate from the typical behavior of the web application and exhibit a different structure compared to regular parameter values. For instance, as shown in Figure 4(b), an invalid DNI (Spanish national ID number) was marked as an anomaly. We divided the dataset into two segments: the training portion, which included 36,000 normal requests and was exclusively utilized for representation learning, and the inference portion, which comprised both 36,000 normal and 25,000 anomalous traffic that was encoded by the model and employed for detection. It has 38 different endpoints, 8 of which have no normal representation and were excluded from our experiment.
@@
The API Traffic Research Dataset Framework (ATRDF) (Lavian et al., 2023) is a recently published HTTP dataset publicly available which includes 18 different API endpoints. The dataset includes attacks such as Directory Traversal, Cookie Injection, Log4j (Figure 4(c)), RCE, Log Forging, SQLi, and XSS. The dataset contains 54,$0$ normal and 78,000 abnormal sets of requests and responses. The training set represents 80% of the normal samples, with the remaining 20% and the abnormal samples used for testing.
@@
In response to the unavailability of a publicly accessible pre-trained model specialized for the API domain, a generic language model was developed to establish a reliable baseline for various detection models. However, developing a robust language model necessitates a significant amount of training data. To address this requirement, an extensive dataset of 1,061,095 API examples was collected. This dataset was obtained by performing a comprehensive data collection process, involving the crawling of a random sample of websites from the Tranco top websites list [^3] https://tranco-list.eu/. The Tranco list is an invaluable resource for cyber-security research as it provides a publicly available compilation of the top one million most popular domains, ranked based on a combination of four reputable lists: Alexa, Cisco Umbrella, Majestic, and Quantcast. Additionally, the Tranco list offers the advantage of being able to filter out unavailable or malicious domains, making it a valuable asset for our research (Pochat et al., 2018).
@@
Figure 4: An example of anomaly requests from CSIC 2010 and ATRDF 2023 datasets where endpoint definition is marked in green, malicious payload is marked in red.
5 Evaluation Metrics
In the realm of anomaly detection, it is commonplace to integrate a binary classification layer into the model architecture. This is due to the fundamental objective of distinguishing between normal and abnormal instances. To evaluate our architecture, we suggest using several performance measures in various experiments, including precision, recall, accuracy, and F1-score. Actual values are represented as True and False by (1) and (0), respectively, and predicted as Positive and Negative values by (1) and (0), respectively. Predicted possibilities of classification models are obtained through the expressions TP, TN, FP, and FN.
@@ 5 Evaluation Metrics
In the current problem, two classes are represented by Positive and Negative, where the positive class corresponds to an abnormal API request and the negative class corresponds to a normal API request. We also compared the training and inference times of the anomaly detection SOTAs and the queries per second (QPS) of the ANN algorithms.
6 Experimental Results
To understand better the two datasets, we use the t-distributed Stochastic Neighbor Embedding (t-SNE) method for dimensionality reduction to graphically depict our high-dimensional datasets. The plots in Figure 5 indicate distinct separation among classes within a reduced dimensional space. This suggests that requests with similarities tend to be grouped, enabling the exploration of a neighborhood for any given sample. Close records from different classes suggest that certain requests are quite similar, causing the embedding to overlap between the two classes. The result for ATRDF 2023 shows no overlap and clear separation between the classes while for CSIC 2010, some records from the two classes were found to be similar. We identified that most of the anomalous requests which overlap with normal requests have no malicious payload and are categorized as unintentional illegal requests.
@@ 6 Experimental Results
Figure 5: Visualization of FastText embedding using t-SNE shows the planar representation of the internal high-dimensional organization of the two classes. Each class is labeled with a unique color for clarity. Dimensionality reduction techniques, such as t-SNE, arrange data points so that those close together represent requests that FastText perceives as having similar patterns.
@@
Then, we compare our method with twelve detecting outlying objects in multivariate data baseline models (Zhao et al., 2019): AE, Feature Bagging (FB), Histogram-based Outlier Detection (HBOS), IF, LOF, Minimum Covariance Determinant (MCD), One-class SVM (OCSVM), Principal Component Analysis (PCA), Deep One-Class Classification for outlier detection (DeepSVDD), Clustering-Based Local Outlier Factor (CBLOF), Kernel Density Estimation (KDE), Linear Model Deviation-based Outlier Detection (LMDD). We apply default hyper-parameters as provided by the original source code across all models for consistency and use the same pre-processed dataset to facilitate optimal comparison.
@@
Furthermore, we conducted an ANN benchmark to effectively contrast our framework with other nearest-neighbor algorithms. This benchmark involved generating algorithm instances based on a configuration file written in YAML format that defines the different methods and algorithms. At the top level, the point type is specified, followed by the distance metric, and finally, each algorithm implementation to be evaluated. Each implementation specifies the Python library, and additional entries provide the necessary arguments. For clarity, an illustrative example of this configuration file is presented in Figure 6, while the complete file is available in the project GitHub repository (Aharon et al., 2024). Both the "space" and "run_groups" lists encompass arguments that should be included at the beginning of every invocation. Each algorithm defines one or more "run_groups", each of which is expanded into several lists of arguments. The Cartesian product of these entries results in numerous argument lists. For instance, consider the Hnswlib entry depicted in Figure  6. This expands into three distinct algorithm instances: Cosine (Cosine Similarity), L2 (Squared L2), and IP (Inner product). Each of these instances undergoes training before being utilized for various experiments. Initially, experiments are conducted with different values of k, representing the number of neighbors to return (e.g., [10, 50, 100, 300, 400, 500, 1000, 2000, 2500, 3000]). Subsequently, experiments are conducted with varying ef_construction values, which denote the size of the dynamic list used during index construction. A larger ef_construction value indicates a higher quality index but also results in longer build times (e.g., [10, 20, 40, 80, 120, 200, 400, 600, 800]). Throughout each run, pertinent information is recorded, including the algorithm name, the time taken to construct the data structure used for indexing, and the outcomes of every query. These query outcomes encompass the neighboring points returned by the algorithm, the duration required to locate these neighbors, and the proximity between the neighbors and the query point. We leveraged the query results to compute the corresponding confusion matrix, enabling us to thoroughly evaluate the classification-by-retrieval performance of each algorithm. We conducted our benchmarking analysis by assessing the performance of various algorithm implementations, including Nmslib, Hnswlib, Bruteforce Blas, BallTree, KDtree, CKDtree, Annoy, Faiss, and RPForest, all of which were evaluated using the publicly accessible ANN-Benchmarks tool as a reference framework (Aumüller et al., 2020). To gain a deeper comprehension of how the embedding layer influences the detection outcome, we assessed all anomaly detection baseline models and the ANN benchmark using two additional prominent language models: BERT (Devlin et al., 2018), and RoBERTa (Liu et al., 2019). Each model was individually trained from scratch and subsequently subjected to evaluation. We employed the RoBERTa model, specifically the RoBERTaForMaskedLM class, with a language modeling head on top [^4] https://huggingface.co/docs/transformers/model_doc/RoBERTa. This model was trained with a maximum sequence length of 512, utilizing 12 hidden layers and 12 attention heads. The training process spanned 10 epochs, employing a batch size of 16, which aligns with a similar approach in a prior study by  (Gniewkowski et al., 2021). Likewise, we employed the BERT model, specifically BertForMaskedLM, which also incorporates a language modeling head on top [^5] https://huggingface.co/docs/transformers/model_doc/bert. This model underwent training with a maximum sequence length of 512, utilizing 4 hidden layers and 4 attention heads, as was the case in a similar task outlined in  (Guo et al., 2021).
@@
We used Intel(R) Xeon(R) Silver 4214R CPU @ 2.40GHz to evaluate the effectiveness of each technique. Each endpoint was evaluated separately, and the average score for all endpoints was used to measure overall detector performance. We evaluate model performance by measuring execution time, inference time, and F1-score during training and testing. It is worth noting that detecting anomalies in real-time is critical in mitigating the effects of an attack. Therefore, it is advisable to analyze each phase separately to better understand each model’s performance. Additionally, we aim to achieve better results compared to previous unsupervised SOTAs. Naturally, supervised-related works should not be compared since they rely on labeled data.
@@
The results for the CSIC 2010 dataset, as can be seen from Table 3, comparing against the conventional outlier detection baselines, FT-ANN and PCA, demonstrate the shortest train times, with 0.0552 and 0.0575 seconds, respectively, whereas AE and LMDD are significantly slower. CBLOF has a relatively fast inference time but does not support incremental learning, requiring access to the entire dataset to fit the model. FT-ANN achieves the highest F1 score of 0.9713, indicating its balanced precision-recall performance. LMDD appears to exhibit the comparatively weakest performance across multiple metrics. This could be attributed to its reliance on a dissimilarity function that may not be well-suited to the complex and diverse anomalies present (Arning et al., 1996).
F1 Score Precision Recall Train Time (Sec) Test Time (Sec)
BERT FastText RoBERTa BERT FastText RoBERTa BERT FastText RoBERTa BERT FastText RoBERTa BERT FastText RoBERTa
FT-ANN 0.9675 0.9713 0.9664 0.947 0.9538 0.9465 0.9982 0.9954 0.9972 0.4458 0.0552 0.5626 0.0353 0.0075 0.0347
MCD 0.9622 0.9572 0.9517 0.9377 0.9382 0.9159 0.9971 0.9774 0.9985 1269.2 6.3603 1261.2 2.4071 0.0021 2.5126
CBLOF 0.9404 0.9445 0.9365 0.9297 0.9318 0.9276 0.9531 0.9589 0.9477 2.8296 0.5747 2.7095 0.0245 0.001 0.0259
IF 0.9397 0.9419 0.9362 0.9289 0.9323 0.9288 0.9525 0.9538 0.9464 16.066 0.5734 15.979 0.3801 0.048 0.4007
HBOS 0.9383 0.9393 0.9367 0.9289 0.9311 0.9274 0.9501 0.9502 0.9483 0.9967 0.8365 1.0126 0.0587 0.0159 0.0613
PCA 0.938 0.9391 0.9349 0.929 0.9304 0.9271 0.9494 0.9504 0.9454 27.829 0.0574 28.412 2.635 0.0025 2.7356
AE 0.938 0.9389 0.9349 0.929 0.9303 0.9271 0.9494 0.9502 0.9454 82.872 75.0644 84.402 0.0922 0.0668 0.0978
KDE 0.937 0.9361 0.9338 0.9283 0.9296 0.9263 0.9483 0.946 0.9441 86.800 3.671 83.525 8.6993 0.3875 8.9307
OCSVM 0.9368 0.936 0.9326 0.9283 0.9295 0.9255 0.948 0.9459 0.9426 45.390 1.6392 50.025 2.8314 0.0848 3.4049
FB 0.9355 0.9354 0.9319 0.9279 0.9303 0.9255 0.946 0.9444 0.9415 10.875 10.5476 10.775 1.199 0.9365 1.2491
LOF 0.9354 0.935 0.9317 0.9279 0.9301 0.9255 0.9459 0.9439 0.9412 1.1774 1.0486 1.1146 0.1291 0.0909 0.1295
DeepSVDD 0.9187 0.9235 0.8999 0.923 0.9381 0.9155 0.921 0.9126 0.8954 38.213 32.7192 38.2421 0.0728 0.0486 0.0679
LMDD 0.3292 0.3017 0.3101 0.5701 0.528 0.5885 0.2476 0.2293 0.225 11882 225.87 11229 517.69 11.642 471.22
Table 3: Performance evaluations of FT-ANN on the CSIC 2010 dataset in comparison to baseline models for detecting outlying objects in multivariate data.